A contractor agrees to build a dam 180ft long and 20ft high, 12ft wide at the bottom and 8ftwide at the top for $9.75 a cubic yard. Find his profit if his costs were $10000?

1
Expert's answer
2021-12-06T09:45:23-0500
"Profit=(9.75)(20)(180)(12)\frac{1}{3^3}-10000\\=\$5600"